9 students are competing in a spelling bee. in how many orders can they finish? In how many ways can they finish 1st-4th
Answer by jim_thompson5910(35256)

There are 9! = 9*8*7*6*5*4*3*2*1 = 362,880 ways to order 9 students (from first to ninth).


If we only care about the first four slots, then there are 9*8*7*6 = 3,024 ways to order 9 students in four slots.